# Optimal Computation in Leaderless and Multi-Leader Disconnected Anonymous Dynamic Networks

### 20 Feb 2024 v5

@@ Appendix A The Subroutine ApproxCount and Its Correctness
In this section we define the subroutine $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ introduced in Section 4.2 and invoked in Listing 2. Its starting point is the algorithm in [24, Section 4.2], with the added difficulty that here we have a strand of leader nodes in the view $\mathcal{V}$ hanging from the first leader node $\tau$ in level $L_{s}$, where the anonymity $a(\tau)$ is an unknown number not greater than $\ell$ (as opposed to $a(\tau)=1$, which is assumed in [24]). The algorithm begins by assuming that $a(\tau)$ is the given parameter $x$, and then it makes deductions on the anonymities of other nodes until it is able to make an estimate $n^{\prime}>0$ on the total number of processes, or report failure in the form of an error code $n^{\prime}\in\{-1,-2,-3\}$. In particular, since the algorithm requires the existence of a long-enough strand hanging from $\tau$, it reports failure if some descendants of $\tau$ (in the relevant levels of $\mathcal{V}$) have more than one child. Another important difficulty that is unique to the multi-leader case is that, even if $\mathcal{V}$ contains a long-enough strand of leader nodes, some nodes in the strand may still be branching in the history tree (that is, the chain of leader nodes is branching, but only one branch appears in $\mathcal{V}$). We will first revisit and generalize [24, Section 4.2] in order to formally state our new subroutine and prove its correctness and running time. We will conclude this section with a proof of Lemma 4.2.
@@
We remark that ApproxCount assumes that the network is $1$-union-connected, as this is sufficient for the main result of Section 4.2 to hold for any $T$-union-connected network (see the proof of Theorem 4.3).
@@
Discrepancy $\delta$. Suppose that ApproxCount is invoked with arguments $\mathcal{V}$, $s$, $x$, $\ell$, where $1\leq x\leq\ell$, and let $\tau$ be the first leader node in level $L_{s}$ of $\mathcal{V}$ (if $\tau$ does not exist, the procedure immediately returns the error code $n^{\prime}=-1$). We define the discrepancy $\delta$ as the ratio $x/a(\tau)$. Clearly, $1/\ell\leq\delta\leq\ell$. Note that, since $a(\tau)$ is not a-priori known by the process executing ApproxCount, then neither is $\delta$.
@@
Conditional anonymity. ApproxCount starts by assuming that the anonymity of $\tau$ is $x$, and makes deductions on other anonymities based on this assumption. Thus, we will distinguish between the actual anonymity of a node $a(v)$ and the conditional anonymity $a^{\prime}(v)=\delta a(v)$ that ApproxCount may compute under the initial assumption that $a^{\prime}(\tau)=x=\delta a(\tau)$.
@@
Guessing conditional anonymities. Let $u$ be a node of a history tree, and assume that the conditional anonymities of all its children $u_{1}$, $u_{2}$, …, $u_{k}$ have been computed: such a node $u$ is called a guesser. If $v$ is not among the children of $u$ but it is at their same level, and the red edge $\{v,u\}$ is present with multiplicity $m\geq 1$, we say that $v$ is guessable by $u$. In this case, we can make a guess $g(v)$ on the conditional anonymity $a^{\prime}(v)$:
@@
Lemma A.1.
@@
where the two sums range over all children of $u$ and $v^{\prime}$, respectively (note that $v=v_{j}$ for some $j$), and $m^{\prime}_{i}$ is the multiplicity of the red edge $\{v_{i},u\}$ (so, $m=m^{\prime}_{j}$). Our lemma now easily follows from the above equation and from the definition of conditional anonymity. ∎
Heavy nodes. As the algorithm in [24], also our subroutine ApproxCount assigns guesses in a well-spread fashion, i.e., in such a way that at most one node per level is assigned a guess.
@@
Our proof is by well-founded induction on $w(v)$. Assume for a contradiction that no descendants of $v$ are heavy. Let $v_{1}$, $v_{2}$, …, $v_{k}$ be the “immediate” descendants of $v$ that have been assigned guesses. That is, for all $1\leq i\leq k$, no internal nodes of the black path with endpoints $v$ and $v_{i}$ have been assigned guesses (observe that $k\geq 1$ because, by assumption, $w(v)>a^{\prime}(v)=\delta a(v)\geq a(v)\geq 1$).
@@
Let $v_{d}$ be the deepest of the $v_{i}$’s, which is unique, since the assignment of guesses is well spread. Note that $v_{d}$ has no siblings at all, otherwise we would have $a(v)>\sum_{i}a(v_{i})$. Due to Lemma A.1, we have $g(v_{d})=a^{\prime}(v_{d})$. Thus,
@@
If $g(v)$ is not an integer, there is nothing to prove. Otherwise, because $v$ is heavy, $g(v)=\lfloor g(v)\rfloor\leq w(v)$. Since $v$ has no heavy descendants, Lemma A.2 implies $w(v)\leq a^{\prime}(v)$. Also, by Lemma A.1, $a^{\prime}(v)\leq g(v)$. We conclude that
@@
When the criterion in Lemma A.3 applies to a node $v$, we say that $v$ has been counted. So, counted nodes are nodes that have been assigned a guess, which was then confirmed to be the correct conditional anonymity.
@@
If $s$ is an isle’s root and $F$ is its set of leaves, we have $a(s)\geq\sum_{v\in F}a(v)$, because $s$ may have some descendants in the history tree $\mathcal{H}$ that do not appear in the view $\mathcal{V}$. This is equivalent to $a^{\prime}(s)\geq\sum_{v\in F}a^{\prime}(v)$. If equality holds, then the isle is said to be complete; in this case, we can easily compute the conditional anonymities of all the internal nodes by adding them up starting from the nodes in $F$ and working our way up to $s$.
Listing 3: The subroutine ApproxCount invoked in Listing 2
@@
7Let $\tau$ be the first leader node in $L_{s}$
8Mark all nodes in $\mathcal{V}$ as not guessed and not counted
9Assign $u:=\tau$; assign $a^{\prime}(u):=x$; mark $u$ as counted
@@ 11   Assign $u:=u^{\prime}$; assign $a^{\prime}(u):=x$; mark $u$ as counted
12While there are guessable levels and a counting cut has not been found
13   Let $v$ be a guessable non-counted node of smallest depth in $\mathcal{V}$
@@ 14   Let $L_{t^{\prime}}$ be the level of $v$; assign $t:=\max\{t,t^{\prime}\}$
15   Assign a guess $g(v)$ to $v$ as in Equation 1; mark $v$ as guessed
@@ 19      If $g(v^{\prime})$ is not an integer, return $(-3,t)$
20      Assign $a^{\prime}(v^{\prime}):=g(v^{\prime})$; mark $v^{\prime}$ as counted and not guessed
@@ 22         For each internal node $w$ of $I$
24            Mark $w$ as counted and not guessed
@@
27   Let $C$ be a counting cut between $L_{s}$ and $L_{t}$
28   Let $n^{\prime}=\sum_{v\in C}a^{\prime}(v)$
29   Let $\ell^{\prime}=\sum_{v\text{ leader node in }C}a^{\prime}(v)$
@@ 30   If $\ell^{\prime}<\ell$, return $(-1,t)$
31   If $\ell^{\prime}>\ell$, return $(-3,t)$
32   Return $(n^{\prime},t)$
@@
Overview of ApproxCount. Our subroutine ApproxCount is found in Listing 3. It repeatedly assigns guesses to nodes based on known conditional anonymities, starting from $\tau$ and its descendants. Eventually some nodes become heavy, and the criterion in Lemma A.3 causes the deepest of them to become counted. In turn, counted nodes eventually form isles; the internal nodes of complete isles are marked as counted, which gives rise to more guessers, and so on. In the end, if a counting cut is created, the algorithm checks whether the conditional anonymities of the leader nodes in the cut add up to $\ell$.
@@
Algorithmic details of ApproxCount. The algorithm ApproxCount uses flags to mark nodes as “guessed” or “counted”; initially, no node is marked. Thanks to these flags, we can check if a node $u\in\mathcal{V}$ is a guesser: let $u_{1}$, $u_{2}$, …, $u_{k}$ be the children of $u$ that are also in $\mathcal{V}$ (recall that a view does not contain all nodes of a history tree); $u$ is a guesser if and only if it is marked as counted, all the $u_{i}$’s are marked as counted, and $a^{\prime}(u)=\sum_{i}a^{\prime}(u_{i})$ (which implies $a(u)=\sum_{i}a(u_{i})$, and thus no children of $u$ are missing from $\mathcal{V}$).
@@
ApproxCount will ensure that nodes marked as guessed are well-spread at all times; if a level of $\mathcal{V}$ contains a guessed node, it is said to be locked. A level $L_{t}$ is guessable if it is not locked and has a non-counted node $v$ that is guessable, i.e., there is a guesser $u$ in $L_{t-1}$ and the red edge $\{v,u\}$ is present in $\mathcal{V}$ with positive multiplicity.
@@
The algorithm starts by assigning a conditional anonymity $a^{\prime}(\tau)=x$ to the first leader node $\tau\in L_{s}$. (If no leader node exists in $L_{s}$, it immediately returns the error code $-1$, Line 6.) It also finds the longest strand $P_{\tau}$ hanging from $\tau$, assigns the same conditional anonymity $x$ to all of its nodes (including the unique child of the last node of $P_{\tau}$) and marks them as counted (Lines 7–11). Then, as long as there are guessable levels and no counting cut has been found yet, the algorithm keeps assigning guesses to non-counted nodes (Line 12).
@@
When a guess is made on a node $v$, some nodes in the path from $v$ to its ancestor in $L_{s}$ may become heavy; if so, let $v^{\prime}$ be the deepest heavy node. If $g(v^{\prime})$ is not an integer, the algorithm returns the error code $-3$ (Line 19). (As we will prove later, this can only happen if $\delta\neq 1$ or some nodes in the strand $P_{\tau}$ have children that are not in the view $\mathcal{V}$.) Otherwise, if $g(v^{\prime})$ is an integer, the algorithm marks $v^{\prime}$ as counted (Line 20), in accordance with Lemma A.3. Furthermore, if the newly counted node $v^{\prime}$ is the root or a leaf of a complete isle $I$, then the conditional anonymities of all the internal nodes of $I$ are determined, and such nodes are marked as counted; this also unlocks their levels if such nodes were marked as guessed (Lines 21–24).
@@
In the end, the algorithm performs a “reality check” and possibly returns an estimate $n^{\prime}$ of $n$, as follows. If no counting cut was found, the algorithm returns the error code $-2$ (Line 25). Otherwise, a counting cut $C$ has been found. The algorithm computes $n^{\prime}$ (respectively, $\ell^{\prime}$) as the sum of the conditional anonymities of all nodes (respectively, all leader nodes) in $C$. If $\ell^{\prime}=\ell$, then the algorithm returns $n^{\prime}$ (Line 32). Otherwise, it returns the error code $-1$ if $\ell^{\prime}<\ell$ (Line 30) or the error code $-3$ if $\ell^{\prime}>\ell$ (Line 31). In all cases, the algorithm also returns the maximum depth $t$ of a guessed or counted node (excluding $\tau$ and its descendants), or $s$ if no such node exists.
@@
Consistency condition. In order for our algorithm to work properly, a condition has to be satisfied whenever a new guess is made. Indeed, note that all of our previous lemmas on guesses rest on the assumption that the conditional anonymities of a guesser and all of its children are known. However, while the node $\tau$ has a known conditional anonymity (by definition, $a^{\prime}(\tau)=x$), the same is not necessarily true of the descendants of $\tau$ and all other nodes that are eventually marked as counted by the algorithm. This justifies the following definition.
@@
During the execution of ApproxCount, if a guess is made on a node $v$ at level $L_{t^{\prime}}$ of $\mathcal{V}$, then $\tau$ has a (unique) descendant $\tau^{\prime}\in L_{t^{\prime}}$ and $a(\tau)=a(\tau^{\prime})$.
@@
Lemma A.4.
Assume that $\delta\geq 1$. Then, as long as 1 is satisfied, the following hold.
@@
Statement (i) is true by design with no additional assumptions, because the algorithm always makes a new guess in a guessable level, which is not locked by definition. Thus, no two nodes marked as guessed are ever in the same level, and so they are well spread.
All other statements can be proved collectively by induction. They certainly hold the first time Line 13 is ever reached. Indeed, the only nodes marked as counted up to this point are $\tau$ and some of its descendants, which are assigned the conditional anonymity $x$. Since $s=t$ and $\tau$ has conditional anonymity $x$ by definition, statement (iv) is satisfied. Note that some descendants of $\tau$ that are marked as counted may not have been assigned their correct conditional anonymities, because some branches of the history tree may not appear in $\mathcal{V}$. However, no guesses have been made yet, and therefore no nodes are heavy; thus, statement (ii) is satisfied. Moreover, the only isles are formed by $\tau$ and its descendants, and are obviously all trivial; so, statement (iii) is satisfied.
@@
The next guess on a new node $v$ is performed properly, as well. Indeed, 1 states that $\tau$ has a descendant $\tau^{\prime}$ at the same level as $v$ such that $a(\tau^{\prime})=a(\tau)$, and therefore $a^{\prime}(\tau^{\prime})=a^{\prime}(\tau)=x$; so, $\tau^{\prime}$ has the correct conditional anonymity. Thus, regardless of what the guesser of $v$ is (either the parent of $\tau^{\prime}$ or some other counted node), the guess at Line 15 is computed properly.
@@
Otherwise, by Lemma A.3, the deepest heavy node $v^{\prime}$ on $P_{v}$ has a correct guess and can be marked as counted, provided that the guess is an integer. Thus, statement (iv) is still true after Line 20. At this point, there are no heavy nodes left, because $v^{\prime}$ is no longer guessed and all of its ancestors along $P_{v}$ end up having the same weight they had before the guess on $v$ was made.
Now, because statement (iii) held before marking $v^{\prime}$ as counted, there can be at most one non-trivial complete isle, and $v^{\prime}$ must be its root or one of its leaves. Note that, due to statement (iv), any isle $I$ identified as complete at Line 21 is indeed complete according to our definitions. Since $I$ is complete, computing the conditional anonymities of its internal nodes as in Line 23 is correct, and therefore statement (iv) is still true after Line 24. Also, the unique non-trivial isle $I$ gets partitioned into non-trivial isles, and statement (iii) holds again. Finally, since Lines 21–24 may only cause weights to decrease, statement (ii) keeps being true. ∎
Running time. We will now study the running time of ApproxCount. We will prove two lemmas that allow us to give an upper bound on the number of rounds it takes for the algorithm to return an output, provided that some conditions are satisfied.
Lemma A.5.
Assume that $\delta\geq 1$. Then, as long as 1 holds, whenever Line 13 is reached, at most $\delta n$ levels are locked.

### 05 Apr 2024 v6

@@ 5.3 Subroutine ApproxCount
Overview. In ApproxCount, we are given a view $\mathcal{V}$ with a strand of leader nodes hanging from the first leader node $\vartheta$ in level $L_{s}$, where the anonymity $a(\vartheta)$ is an unknown number not greater than $\ell$. The algorithm begins by assuming that $a(\vartheta)$ is the given parameter $x$, and then it makes deductions on the anonymities of other nodes until it is able to make an estimate $n^{\prime}>0$ on the total number of processes, or report failure in the form of an error code $n^{\prime}\in\{-1,-2,-3\}$. In particular, since the algorithm requires the existence of a long-enough strand hanging from $\vartheta$, it reports failure if some descendants of $\vartheta$ (in the relevant levels of $\mathcal{V}$) have more than one child.
An important difficulty that is unique to the multi-leader case is that, even if $\mathcal{V}$ contains a long-enough strand of leader nodes, some nodes in the strand may still be branching in the history tree, and all branches except one may be missing from $\mathcal{V}$.
@@
We remark that ApproxCount assumes that the network is 1-union-connected, as this is sufficient for the main result of Section 5.2 to hold for any $\tau$-union-connected network (see the proof of Theorem 5.2).
@@
Discrepancy $\delta$. Suppose that ApproxCount is invoked with arguments $\mathcal{V}$, $s$, $x$, $\ell$, where $1\leq x\leq\ell$, and let $\vartheta$ be the first leader node in level $L_{s}$ of $\mathcal{V}$ (if $\vartheta$ does not exist, the procedure immediately returns the error code $n^{\prime}=-1$). We define the discrepancy $\delta$ as the ratio $x/a(\vartheta)$. Clearly, $1/\ell\leq\delta\leq\ell$. Note that, since $a(\vartheta)$ is not a-priori known by the process executing ApproxCount, then neither is $\delta$.
@@
Conditional anonymity. ApproxCount starts by assuming that the anonymity of $\vartheta$ is $x$, and makes deductions on other anonymities based on this assumption. Thus, we will distinguish between the actual anonymity of a node $a(v)$ and the conditional anonymity $a^{\prime}(v)=\delta a(v)$ that ApproxCount may compute under the initial assumption that $a^{\prime}(\vartheta)=x=\delta a(\vartheta)$.
@@
Guessing conditional anonymities. Let $u$ be a node of a history tree, and assume that the conditional anonymities of all its children $u_{1}$, $u_{2}$, …, $u_{k}$ have been computed: such a node $u$ is called a guesser. If $v$ is not among the children of $u$ but is at their same level, and the red edge $\{v,u\}$ is present with multiplicity $m\geq 1$, we say that $v$ is guessable by $u$ (see Figure 6). In this case, we can make a guess $g(v)$ on the conditional anonymity $a^{\prime}(v)$:
@@
Figure 6: If the anonymities of $u$, $u_{1}$, $u_{2}$, …, $u_{k}$ are known, then $v$ is guessable by $u$.
Lemma 5.3.
@@
where the two sums range over all children of $u$ and $v^{\prime}$, respectively (note that $v=v_{j}$ for some $j$), and $m^{\prime}_{i}$ is the multiplicity of the red edge $\{v_{i},u\}$ (so, $m=m^{\prime}_{j}$). Therefore, we have the inequality
$a(v)\leq\frac{\sum_{i}a(u_{i})\,m_{i}}{m}$
which becomes an equality if $v$ has no siblings. Thus,
$a^{\prime}(v)=\delta a(v)\leq\frac{\sum_{i}\delta a(u_{i})\,m_{i}}{m}=\frac{\sum_{i}a^{\prime}(u_{i})\,m_{i}}{m}=g(v)$
and so $g(v)\geq a^{\prime}(v)$, with equality if $v$ has no siblings. ∎
Heavy nodes. The subroutine ApproxCount assigns guesses in a well-spread fashion, that is, in such a way that no two sibling nodes are assigned a guess. In other words, at most one of the children of each node is assigned a guess.
@@
Our proof is by well-founded induction on $w(v)$. Assume for a contradiction that no descendants of $v$ are heavy. Let $v_{1}$, $v_{2}$, …, $v_{k}$ be the “immediate” descendants of $v$ that have been assigned guesses. That is, for all $1\leq i\leq k$, no internal nodes of the black path with endpoints $v$ and $v_{i}$ have been assigned guesses. Observe that $k\geq 1$ because, by assumption, $w(v)>a^{\prime}(v)=\delta a(v)\geq a(v)\geq 1$.
@@
Let $1\leq d\leq k$ be such that $v_{d}$ has maximum depth. Since the assignment of guesses is well spread, no sibling of $v_{d}$ has been assigned a guess. However, since $a(v)=\sum_{i}a(v_{i})$, it follows that $v_{d}$ has no siblings at all, for otherwise $a(v)>\sum_{i}a(v_{i})$. Due to Lemma 5.3, we have $g(v_{d})=a^{\prime}(v_{d})$. Thus,
@@
If $g(v)$ is not an integer, there is nothing to prove. Otherwise, because $v$ is heavy, $g(v)=\lfloor g(v)\rfloor\leq w(v)$. Since $v$ has no heavy descendants, Lemma 5.4 implies $w(v)\leq a^{\prime}(v)$. Also, by Lemma 5.3, $a^{\prime}(v)\leq g(v)$. We conclude that
@@
When the criterion in Lemma 5.5 applies to a node $v$, we say that $v$ has been counted. So, counted nodes are nodes that have been assigned a guess, which was then confirmed to be the correct conditional anonymity.
@@
If $s$ is an isle’s root and $F$ is its set of leaves, we have $a(s)\geq\sum_{v\in F}a(v)$, which is equivalent to $a^{\prime}(s)\geq\sum_{v\in F}a^{\prime}(v)$. Note that equality does not necessarily hold, because $s$ may have some descendants in the history tree $\mathcal{H}$ that do not appear in the view $\mathcal{V}$. If equality holds, then the isle is said to be complete. In this case, given the conditional anonymities of $s$ and of all nodes in $F$, we can easily compute the conditional anonymities of all the internal nodes by adding them up starting from the nodes in $F$ and working our way up to $s$.
Figure 7: The first levels of (a view of) a history tree, where the colored nodes are counted. The blue nodes form a counting cut, and the orange ones define a non-trivial isle with root $s$, where the nodes with a dot are internal.
Listing 3: The subroutine ApproxCount invoked in Listing 2.
@@
7Let $\vartheta$ be the first leader node in $L_{s}$
8Mark all nodes in $\mathcal{V}$ as not guessed, not counted, and not locked
9Assign $u:=\vartheta$; assign $a^{\prime}(u):=x$; mark $u$ as counted
@@ 11   Assign $u:=u^{\prime}$; assign $a^{\prime}(u):=x$; mark $u$ as counted
12While there are guessable, non-counted, non-locked nodes in $\mathcal{V}$ and a counting cut has not been found
13   Let $v$ be a guessable, non-counted, non-locked node of smallest depth in $\mathcal{V}$
@@ 14   Let $L_{t^{\prime}}$ be the level of $v$; assign $t:=\max\{t,t^{\prime}\}$
15   Assign a guess $g(v)$ to $v$ as in Equation 1; mark $v$ as guessed; lock $v$ and all its siblings
@@ 19      If $g(v^{\prime})$ is not an integer, return $(-3,t)$
20      Assign $a^{\prime}(v^{\prime}):=g(v^{\prime})$; mark $v^{\prime}$ as counted and not guessed; unlock $v^{\prime}$ and all its siblings
@@ 22         For each internal node $w$ of $I$
24            Mark $w$ as counted, not guessed, and not locked
@@
27   Let $C$ be a counting cut (between $L_{s}$ and $L_{t}$)
28   Assign $n^{\prime}:=\sum_{v\in C}a^{\prime}(v)$
29   Assign $\ell^{\prime}:=\sum_{v\text{ leader node in }C}a^{\prime}(v)$
@@ 30   If $\ell^{\prime}<\ell$, return $(-1,t)$
31   Else if $\ell^{\prime}>\ell$, return $(-3,t)$
32   Else return $(n^{\prime},t)$
@@
Overview of ApproxCount. Our subroutine ApproxCount is found in Listing 3. It repeatedly assigns guesses to nodes based on known conditional anonymities, starting from $\vartheta$ and its descendants. Eventually some nodes become heavy, and the criterion in Lemma 5.5 causes the deepest of them to become counted. In turn, counted nodes eventually form isles; the internal nodes of complete isles are marked as counted, which gives rise to more guessers, and so on. In the end, if a counting cut is created, the algorithm checks whether the conditional anonymities of the leader nodes in the cut add up to $\ell$.
@@
Algorithmic details of ApproxCount. The algorithm ApproxCount uses flags to mark nodes as “guessed”, “counted”, or “locked”; initially, no node is marked. Thanks to these flags, we can check if a node $u\in\mathcal{V}$ is a guesser: let $u_{1}$, $u_{2}$, …, $u_{k}$ be the children of $u$ that are also in $\mathcal{V}$ (recall that a view does not contain all nodes of a history tree); $u$ is a guesser if and only if it is marked as counted, all the $u_{i}$’s are marked as counted, and $a^{\prime}(u)=\sum_{i}a^{\prime}(u_{i})$ (which implies $a(u)=\sum_{i}a(u_{i})$, and thus no children of $u$ are missing from $\mathcal{V}$).
@@
ApproxCount will ensure that nodes marked as guessed are well-spread at all times: if a node of $\mathcal{V}$ is guessed, all its siblings (including itself) become locked. While a node is locked, it cannot receive guesses. As defined earlier, a node $v$ in level $L_{t}$ of $\mathcal{V}$ is guessable if there is a guesser $u$ in $L_{t-1}$ and the red edge $\{v,u\}$ is present in $\mathcal{V}$ with positive multiplicity.
@@
The algorithm starts by assigning a conditional anonymity $a^{\prime}(\vartheta)=x$ to the first leader node $\vartheta\in L_{s}$. If no leader node exists in $L_{s}$, the error code $-1$ is immediately returned (Line 6). The algorithm also finds the longest strand $P_{\vartheta}$ hanging from $\vartheta$, assigns the same conditional anonymity $x$ to all of its nodes (including the unique child of the last node of $P_{\vartheta}$) and marks them as counted (Lines 7–11). Then, as long as there are nodes that can receive a guess and no counting cut has been found, the algorithm keeps assigning guesses to nodes. A node can receive a guess if it is guessable, not counted, and not locked (Line 12).
@@
When a guess is made on a node $v$, the node itself and all of its siblings become locked (Line 15). This is to ensure that guessed nodes will always be well spread. Moreover, as a result of $v$ becoming guessed, some nodes in the path from $v$ to its ancestor in $L_{s}$ may become heavy; if this happens, let $v^{\prime}$ be the deepest heavy node (Line 18). If $g(v^{\prime})$ is not an integer, the algorithm returns the error code $-3$ (Line 19). (As we will prove later, this can only happen if $\delta\neq 1$ or some nodes in the strand $P_{\vartheta}$ have children that are not in the view $\mathcal{V}$.) Otherwise, if $g(v^{\prime})$ is an integer, the algorithm marks $v^{\prime}$ as counted and not guessed, in accordance with Lemma 5.5. Also, since $v^{\prime}$ is no longer guessed, its siblings become unlocked and are again eligible to receive guesses (Line 20). Furthermore, if the newly counted node $v^{\prime}$ is either the root or a leaf of a complete isle $I$, then the conditional anonymities of all the internal nodes of $I$ are determined, and such nodes are marked as counted (Lines 21–24).
@@
In the end, the algorithm performs a “reality check” and possibly returns an estimate $n^{\prime}$ of $n$ as follows. If no counting cut was found, the algorithm returns the error code $-2$ (Line 25). Otherwise, a counting cut $C$ has been found. The algorithm computes $n^{\prime}$ (respectively, $\ell^{\prime}$) as the sum of the conditional anonymities of all nodes (respectively, all leader nodes) in $C$. If $\ell^{\prime}=\ell$, then the algorithm returns $n^{\prime}$ (Line 32). Otherwise, it returns the error code $-1$ if $\ell^{\prime}<\ell$ (Line 30) or the error code $-3$ if $\ell^{\prime}>\ell$ (Line 31). In all cases, the algorithm also returns the maximum depth $t$ of a guessed or counted node (excluding $\vartheta$ and its descendants), or $s$ if no such node exists.
@@
Consistency condition. In order for our algorithm to work properly, a condition has to be satisfied whenever a new guess is made. Indeed, note that all of our previous lemmas on guesses rest on the assumption that the conditional anonymities of a guesser and all of its children are known. However, while the node $\vartheta$ has a known conditional anonymity (by definition, $a^{\prime}(\vartheta)=x$), the same is not necessarily true of the descendants of $\vartheta$ and all other nodes that are eventually marked as counted by the algorithm. This justifies the following definition.
@@
During the execution of ApproxCount, if a guess is made on a node $v$ at level $L_{t^{\prime}}$ of $\mathcal{V}$, then $\vartheta$ has a (unique) descendant $\vartheta^{\prime}\in L_{t^{\prime}}$ and $a(\vartheta)=a(\vartheta^{\prime})$.
@@
Lemma 5.6.
Assume that $\delta\geq 1$. Then, as long as 1 is satisfied, the following statements hold.
@@
Statement (i) is true by design with no additional assumptions, because the algorithm only makes new guesses on unlocked nodes. In turn, a node is locked if and only if it is marked as guessed or has a sibling that is marked as guessed. Thus, no two nodes marked as guessed can be siblings, e.g., guesses are well spread.
All other statements can be proved collectively by induction. They certainly hold the first time Line 13 is ever reached. Indeed, the only nodes marked as counted up to this point are $\vartheta$ and some of its descendants, which are assigned the conditional anonymity $x$. Since $s=t$ and $\vartheta$ has conditional anonymity $x$ by definition, statement (iv) is satisfied. Note that some descendants of $\vartheta$ that are marked as counted may not have been assigned their correct conditional anonymities, because some branches of the history tree may not appear in $\mathcal{V}$. However, no guesses have been made yet, and therefore no nodes are heavy; thus, statement (ii) is satisfied. Moreover, the only isles are formed by $\vartheta$ and its descendants, and are obviously all trivial; so, statement (iii) is satisfied.
@@
The next guess on a new node $v$ is performed properly, as well. Indeed, 1 states that $\vartheta$ has a descendant $\vartheta^{\prime}$ at the same level as $v$ such that $a(\vartheta^{\prime})=a(\vartheta)$, and therefore $a^{\prime}(\vartheta^{\prime})=a^{\prime}(\vartheta)=x$; so, $\vartheta^{\prime}$ has the correct conditional anonymity. Thus, regardless of what the guesser of $v$ is (either the parent of $\vartheta^{\prime}$ or some other counted node), the guess at Line 15 is computed properly.
@@
Otherwise, by Lemma 5.5, the deepest heavy node $v^{\prime}$ on $P_{v}$ has a correct guess and can be marked as counted, provided that the guess is an integer. Thus, statement (iv) is still true after Line 20. At this point, there are no heavy nodes left, because $v^{\prime}$ is no longer guessed and all of its ancestors along $P_{v}$ end up having the same weight they had before the guess on $v$ was made.
Now, because statement (iii) held before marking $v^{\prime}$ as counted, there can be at most one non-trivial complete isle, and $v^{\prime}$ must be its root or one of its leaves. Note that, due to statement (iv), any isle $I$ identified as complete at Line 21 is indeed complete according to our definitions. Since $I$ is complete, computing the conditional anonymities of its internal nodes as in Line 23 is correct, and therefore statement (iv) is still true after Line 24. Also, the unique non-trivial isle $I$ is reduced to trivial isles, and statement (iii) holds again. Finally, since Lines 21–24 may only cause weights to decrease, statement (ii) keeps being true. ∎
Running time. We will now study the running time of ApproxCount. We will prove two lemmas that allow us to give an upper bound on the number of rounds it takes for the algorithm to return an output.
Recall that a node $v$ of the history tree $\mathcal{H}$ is said to be missing from level $L_{i}$ of the view $\mathcal{V}$ if $v$ is at the level of $\mathcal{H}$ corresponding to $L_{i}$ but does not appear in $\mathcal{V}$. Clearly, if a level of $\mathcal{V}$ has no missing nodes, all previous levels have no missing nodes, either.
Lemma 5.7.
Assume that $\delta\geq 1$. Then, as long as 1 holds, whenever Line 13 is reached, at most $\delta(n-1)$ levels contain locked nodes.